if points A,Band C on a coordinate line have coordinates -5,3 and 10 respectively,find d(A,C)+d(B,C).thank you.
Answer by venugopalramana(3286)

if points A,Band C on a coordinate line have coordinates -5,3 and 10 respectively,find d(A,C)+d(B,C).thank you.
DO YOU MEAN A,B,C HAVE SAME X (OR Y ) COORDINATE? IF SO
AB = 3-(-5)=3+5=8
BC = 10-3=7
AC=10+5=15
AC+BC=7+15=22